OpenStudy (anonymous):

find the area and perimeter of a square with a diagonal of 10 inches please help :) thanks

OpenStudy (anonymous):

\[x^{2}+x ^{2}=10^{2}\]\[2x^{2}=100\]\[x^{2}=\frac{100}{2}\]\[x=\sqrt{50}\]\[A=\sqrt{50}\times \sqrt{50} \] A=50 \[P=4\sqrt{50}\] P=28.3

OpenStudy (anonymous):

the measure of its side is \[5\sqrt{2}\] so, we are going to use the formula in finding its perimeter which is A=4s =4(5\[4(5\sqrt{2}\]) its perimeter is \[20\sqrt{2}\] inches \[A=s ^{2}\] \[A=(5\sqrt{2})^{2}\] \[A=25(2)\] A=50 in^2